OpenStudy (riot2015):

Find the simplified form of the expression. Give your answer in scientific notation. (5*10^6)(3*10^5) A) 1.5*10^11 B) 1.5*10^13 C) 15*10^11 D) 1.5*10^10

OpenStudy (unklerhaukus):

\[(5\times10^6)(3\times10^5)\\ = (5\times3)\times10^{6+5}\\ = \] what do you get?

OpenStudy (riot2015):

C

OpenStudy (unklerhaukus):

NB: a number in scientific notation has only digit before the decimal place.

OpenStudy (riot2015):

A

OpenStudy (unklerhaukus):

\[(5\times10^6)(3\times10^5)\\ = (5\times3)\times10^{6+5}\\ = 15\times10^{11}\\ = \frac{15}{10}\times10^{11+1}\\ = %1.5\times10^{12} \]

OpenStudy (riot2015):

so it's not c because it would be two numbers before the decimal

OpenStudy (riot2015):

Right?

OpenStudy (unklerhaukus):

right, (C is equal to the product, but it is not correct scientific notation)

OpenStudy (riot2015):

B because of 12?

OpenStudy (unklerhaukus):

where is the 12?

OpenStudy (riot2015):

The exponent you put sais 11+1

OpenStudy (riot2015):

This is so hard haha

OpenStudy (unklerhaukus):

i dont see a ...10^12 option in the options above

OpenStudy (riot2015):

Ohhh sorry that 13 is supposed to be 12

OpenStudy (unklerhaukus):

that makes sense then.

OpenStudy (riot2015):

Ok can u help me on one more please?

OpenStudy (unklerhaukus):

Sure.

OpenStudy (riot2015):

(4*10^10)(9*10^-5)

OpenStudy (riot2015):

Same question as the first one

OpenStudy (unklerhaukus):

\[(4\times10^{10})(9\times10^{-5})\\ = (4\times9)\times 10^{10-5}\\ = \] what do you get this time?

OpenStudy (riot2015):

36*10^5?

OpenStudy (unklerhaukus):

yeah, now just get it to be a single digit before the decimal point

OpenStudy (unklerhaukus):

(divide 36 by 10) and (multiply 10^5 by 10)

OpenStudy (unklerhaukus):

what does it become?

OpenStudy (riot2015):

3.6*10

OpenStudy (unklerhaukus):

3.6*10^?

OpenStudy (riot2015):

But when I multiply I get 1000 and that's not one of my choices

OpenStudy (unklerhaukus):

10^5 * 10 = 10^(5+1)

OpenStudy (riot2015):

ohhh so 3.6*10^6

OpenStudy (unklerhaukus):

\[\color{red}\checkmark\]

OpenStudy (unklerhaukus):

great work